# Indecomposable set-theoretical solutions to the Yang–Baxter equation of size $p^{2}$

### v1

@@ Abstract.
The quantum Yang–Baxter equation is a braiding condition on complex vector spaces which is of high relevance in several fields of mathematics, such as knot theory and quantum group theory. A combinatorial approach is the investigation of set-theoretic solutions to the Yang–Baxter equation and their associated algebraic structures. In this article, we focus on indecomposable set-theoretic solutions to the Yang–Baxter equation. More specifically, we give a full classification of those which are of size $p^{2}$.
@@ Introduction
The quantum Yang–Baxter equation, named after the physicists Chen-Ning Yang and Rodney James Baxter, first appeared in the context of statistical mechanics and integrable systems [3, 21]. Given a dimensional complex vector space $V$, call a linear operator $R:V^{\otimes 2}\to V^{\otimes 2}$ a solution to the Yang–Baxter equation if $R$ is bijective and satisfies the following equation within $\mathrm{End}(V^{\otimes 3})$:
@@
Outside of physics, solutions to this equation have been recognized to be of importance in low-dimensional topology, as results of Turaev [20] show that each Yang–Baxter operator gives rise to an associated knot invariant. While these operators are therefore of high importance in knot theory, they are very difficult to construct; initially, the only source of solutions was the theory of quantum groups. However, Drinfeld [8] observed that a particular class of solutions provides a handier way to obtain classes of solutions by demanding that $R$ permutes the tensors $v_{i}\otimes v_{j}$ where $(v_{i})_{i\in I}$ is a basis of $V$. By doing so, one gets a discretized version of the Yang–Baxter equation, the set-theoretic Yang–Baxter equation, which consists of a set $X$ and a bijection $r:X^{2}\to X^{2}$ that satisfies the following equation in the symmetric group $\operatorname{Sym}_{X^{3}}$:
@@
Two set-theoretic solutions $(X,r)$ and $(Y,s)$ are called equivalent if there is a bijection $f:X\to Y$ such that $(f\times f)r=s(f\times f)$.
@@
By the size of a set-theoretic solution, we mean the size of $X$. From now on, when talking about solutions, we mean a set-theoretic solution to the Yang–Baxter equation that is finite, involutive and non-degenerate.
Rump has shown in [17] that the binary operation $x\cdot y=\lambda_{x}^{-1}(y)$ turns $X$ into a particularly easy algebraic structure; a (non-degenerate) cycle set, which is defined as a set $X$ with a binary operation $(x,y)\mapsto x\cdot y$ such that the following axioms are satisfied:
@@
Homomorphisms of cycle sets are defined in the obvious way. Consequently, this holds for all derived notions, such as isomorphisms. Note that isomorphism of cycle sets translates to equivalence of set-theoretic solutions under the correspondence described above. As usual, if there is no danger of confusion, a cycle set $(X,\cdot)$ is simply referred to by its underlying set $X$.
The cycle sets considered in this work will always be assumed to be of finite size and will be referred to as cycle sets, without the predicate nondegenerate. Cycle sets are in bijective correspondence with finite nondegenerate involutive solutions to the Yang-Baxter equation. More precisely, given a cycle set $(X,\cdot)$ then $(X,r_{X})$ is a nondegenerate involutive solution to the YBE, where $r_{X}:X^{2}\to X^{2}$ is given by $r_{X}(x,y)=(\sigma_{x}^{-1}(y),\sigma_{x}^{-1}(y)\cdot x)$.
Of particular importance for understanding general cycle sets is the class of indecomposable cycle sets. A cycle set $X$ is called indecomposable if there are no subsets $X_{1},X_{2}\neq\emptyset$ closed under the cycle set operation such that $X=X_{1}\cup X_{2}$. In a sense, they are elementary building blocks for cycle sets and turn out to be significantly more well-behaved. This can be seen from computations of Akgün, Mereb and Vendramin [1] that imply that the number of cycle sets of given cardinality grows in an aggressive manner, while the number of indecomposable cycle sets stays quite small. This is a hint that there are strong obstructions for a cycle set to be indecomposable.
Indeed, up to isomorphism, there is only one indecomposable cycle set of size $p$, $p$ a prime number, that is given by $X=\mathbb{Z}_{p}$ with the operation $x\cdot y=y+1$ [9, Theorem 2.13]. Furthermore, Jedliĉka and Pilitowska [14] have recently classified all indecomposable cycle sets of multipermutation level 2 by group-theoretic means. In particular, this includes all indecomposable cycle sets of order $pq$ that are of finite multipermutation level, where $p,q$ are (not necessarily distinct) primes. Cedó and Okniński [6, Section 5] have described a class of simple cycle sets of size $p^{2}$, $p$ a prime. In [7], the same authors have recently proven that all indecomposable cycle sets of squarefree size are of finite multipermutation level.
In this article, we build on these results by providing a full classification of indecomposable cycle sets of size $p^{2}$, $p$ a prime.
Note that a cycle set $(X,\cdot)$ is indecomposable if and only if the corresponding set-theoretic solution $(X,r)$ is indecomposable in the sense that there is no partition $X=A\sqcup B$ ($A,B\neq\emptyset$) such that $r(X^{2})=X^{2}$ and $r(Y^{2})=Y^{2}$.
Using the correspondence between cycle sets and set-theoretic solutions, our final result, 5.1, reformulated in terms of set-theoretic solutions, will be as follows:
@@
Each indecomposable solution $(X,r)$ to the Yang–Baxter equation of size $p^{2}$ for some prime $p$ is equivalent to one of the following solutions:
@@
    $r\begin{pmatrix}(a,x)\\ (b,y)\end{pmatrix}=\begin{pmatrix}(b-1,y-S\chi_{0}(b-1)-\Phi(b-1-a))\\ (a+1,x+S\chi_{0}(a)+\Phi(a-b+1))\end{pmatrix}$
    where $\Phi:\mathbb{Z}_{p}\to\mathbb{Z}_{p}$ is a nonconstant map with $\Phi(0)=0$, $S\in\mathbb{Z}_{p}$ and
    $\chi_{0}:\mathbb{Z}_{p}$ $\to\mathbb{Z}_{p}$
@@
    The parameters $S,\Phi$ and $S^{\prime},\Phi^{\prime}$ define equivalent solutions if and only if $S=S^{\prime}$ and $\alpha\Phi=\Phi^{\prime}$ for some $\alpha\in\mathbb{Z}_{p}^{\ast}$.
@@
    where $\Phi:\mathbb{Z}_{p}\to\mathbb{Z}_{p}$ is a nonconstant map with $\Phi(x)=\Phi(-x)$ and $\alpha\in\mathbb{Z}_{p}^{\ast}$ is such that $\Phi(\alpha x)=\alpha\Phi(x)$.
    The parameters $\alpha,\Phi$ and $\alpha^{\prime},\Phi^{\prime}$ define equivalent cycle sets if and only if $\alpha=\alpha^{\prime}$ and there is a $\beta\in\mathbb{Z}_{p}^{\ast}$ such that $\beta^{-1}\Phi(\beta x)=\Phi^{\prime}(x)$ for all $x\in\mathbb{Z}_{p}$.
Note that the classification of indecomposable retractable solutions of size $p^{2}$ follows from the theoretical classification of indecomposable solutions of multipermutation level $2$ in [14]. We merely give them in a more explicit form building further upon their results, which also gives a more uniform presentation compared to the irretractable case. This more explicit form is obtained in Section 2, where we also present a relation between the size of any indecomposable cycle set of finite multipermutation level and the order of its permutation group.
@@ 1. Preliminaries
which is easily verified to be left distributive. The set of fixed points of the $\lambda$-action is denoted
$\operatorname{Fix}(A)=\{a\in A\colon b+a=b\circ a\text{ for all }b\in A\}=\{a\in A\mid b*a=0\text{ for all }b\in A\}$
@@
Let $b\in B$. We have to show that $\lambda_{a}(b)\in B$ for all $a\in A$. Note that ${}^{a}b\in B$ because $B$ is normal, therefore $\lambda_{{}^{a}b}(a)\in a+B$ as $B^{\circ}$ acts trivially on $A^{+}/B^{+}$. Using 1.1 and the normality of $B^{+}$, we deduce that
@@
Let $A$ be a finite skew brace and $B$ a subbrace such that $B^{+}$ is a normal subgroup of $A^{+}$ and $p=|A|/|B|$ is equal to the least prime dividing $|A|$. Then $B$ is an ideal of $A$.
@@
$\operatorname{Fix}(A)$ is the set of fixed points under the action $\lambda:A^{\circ}\to\operatorname{Aut}(A^{+})$. It is well-known that an action of a $p$-group by automorphisms of a $p$-group always has nontrivial fixed points.
If we assume that $A^{\circ}$ is abelian, 1.2 implies that $\operatorname{Fix}(A)\subseteq\operatorname{Soc}(\mathcal{G}),$ hence $\operatorname{Soc}(A)\neq 0$ as well. ∎
Let $G$ be a group acting on a set $X$. A system of imprimitivity is a partition $X=\bigsqcup_{i\in I}X_{i}$ that is invariant under the group action in the sense that for any $g\in G$, $i\in I$, there is a $j\in I$ such that $g\cdot X_{i}=X_{j}$. The subsets $X_{i}$ are called blocks of the system. A system of imprimitivity is trivial if either $|I|=1$ or $|X_{i}|=1$ for all $i\in I$. Else it is nontrivial.
@@
Let $G$ be a nonabelian $p$-group with a transitive action on a set $X$ of size $p^{2}$. Then $G$ has at most one nontrivial system of imprimitivity.
@@
If there is more than one nontrivial system of imprimitivity, a theorem of Lucchini [16, Theorem 1] implies that $G\leqslant\operatorname{Sym}_{p}\times\operatorname{Sym}_{p}$. But if $G$ is a $p$-group, this forces $G$ to be isomorphic to a subgroup of $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ and therefore to be abelian. ∎
Let the groups $G,H$ act on sets $X$ resp. $Y$. Then the wreath product is defined as the semidirect product $G\wr H=G\ltimes H^{X}$ where $G$ acts on $H^{X}$ by $g\cdot(h_{x})_{x\in X}=(h_{g^{-1}\cdot x})_{x\in X}$.
$G\wr H$ acts on $X\times Y$ by permutations of the form
@@
As a special case of [13, II, 3.2 Satz], we know that every solvable transitive subgroup of $\operatorname{Sym}_{\mathbb{Z}_{p}}$ is conjugate to a subgroup of $\mathrm{AGL}(1,p)$. So we conclude that $G$ is conjugate to a subgroup of $\mathrm{AGL}(1,p)\wr\mathrm{AGL}(1,p)$. As $\mathbb{Z}_{p}\wr\mathbb{Z}_{p}$ is a $p$-Sylow subgroup of $\mathrm{AGL}(1,p)\wr\mathrm{AGL}(1,p)$ also the last part of the statement follows. ∎
@@
To a cycle set $(X,\cdot)$ we associate the group
@@
By defining $x+y=x\circ\sigma_{x}(y)$, we then obtain a brace $(G(X),+,\circ)$. The canonical map $\iota:(X,\cdot)\to(G(X),\cdot):x\mapsto x$ is always an injective cycle set homomorphism, so we can identify $X$ with $\iota(X)$. Note that it follows immediately from the definition that $\iota(X)$ is a cycle base of $G(X)$.
@@
Recall that a cycle set $X$ is indecomposable if and only if $\mathcal{G}(X)$ acts transitively on $X$. The above group structure will also be denoted by $(\mathcal{G}(X),\circ)$ and we obtain, as described in [5], a brace structure $(\mathcal{G}(X),+,\circ)$ by defining $\sigma_{x}^{-1}+\sigma_{y}^{-1}=\sigma_{x}^{-1}\circ\sigma^{-1}_{\sigma_{x}(y)}$. Note that this means that we obtain a surjective brace homomorphism $G(X)\to\mathcal{G}(X):x\mapsto\sigma_{x}^{-1}$, where the kernel is precisely $\operatorname{Soc}(G(X))$. As a result, the canonical map
@@
becomes a cycle set homomorphism. The image $\pi(X)$ is called the retraction of $X$, denoted $\mathrm{Ret}(X)$. If $\pi$ is injective, hence $(X,\cdot)$ is isomorphic to a sub-cycle set of $(\mathcal{G}(X),\cdot)$, then we say that $X$ is irretractable. When $\pi$ is not injective, $(X,\cdot)$ is retractable. Note that $\pi(x)=\pi(y)$ if and only if $\sigma_{x}=\sigma_{y}$. If moreover, we can obtain a one-element cycle set by iteratively repeating the retraction process on a cycle set $X$, we say that $X$ has finite multipermutation level.
@@
Let $f:X\twoheadrightarrow Y$ be a surjective cycle set homomorphism. If $X$ is indecomposable, then $Y$ is also indecomposable and the fibres $f^{-1}(y)$ all have the same cardinality. In particular, $|Y|$ divides $|X|$.
2. Indecomposable cycle sets of finite multipermutation level

### v2

# Indecomposable involutive set-theoretical solutions to the Yang–Baxter equation of size $p^{2}$

@@ Abstract.
The quantum Yang–Baxter equation is a braiding condition on vector spaces which is of high relevance in several fields of mathematics, such as knot theory and quantum group theory. Their combinatorial counterpart are set-theoretic solutions to the Yang–Baxter equation, whose investigation is strongly driven by the study of algebraic objects called (skew) braces. In this article, we focus on indecomposable involutive non-degenerate set-theoretic solutions to the Yang–Baxter equation. More specifically, through a thorough analysis of their associated braces, we give a full classification of those which are of size $p^{2}$, for $p$ a prime.
@@ Introduction
The quantum Yang–Baxter equation, named after the physicists Chen-Ning Yang and Rodney James Baxter, first appeared in the context of statistical mechanics and integrable systems [3, 22]. Given a (typically complex) vector space $V$, call a linear operator $R:V^{\otimes 2}\to V^{\otimes 2}$ a solution to the Yang–Baxter equation if $R$ is bijective and satisfies the following equation within $\mathrm{End}(V^{\otimes 3})$:
@@
Outside of physics, solutions to this equation have been recognized to be of importance in low-dimensional topology, as results of Turaev [21] show that each Yang–Baxter operator gives rise to an associated knot invariant. While these operators are therefore of high importance in knot theory, they are very difficult to construct; initially, the only source of solutions was the theory of quantum groups. However, Drinfeld [8] observed that a particular class of solutions provides a handier way to obtain classes of solutions by demanding that $R$ permutes the tensors $v_{i}\otimes v_{j}$ where $(v_{i})_{i\in I}$ is a basis of $V$. By doing so, one gets a discretized version of the Yang–Baxter equation, the set-theoretic Yang–Baxter equation, which consists of a set $X$ and a bijection $r:X^{2}\to X^{2}$ that satisfies the following equation in $\operatorname{Sym}_{X^{3}}$, the symmetric group on the set $X^{3}$:
@@
A homomorphism of cycle sets $f:(X,r)\to(Y,s)$ is a map $f:X\to Y$ such that $(f\times f)r=s(f\times f)$. An isomorphism of cycle sets is a bijective homomorphism.
@@
By the size of a set-theoretic solution, we mean the size of $X$. From now on, when talking about set-theoretic solutions to the Yang–Baxter equation that are finite, involutive and non-degenerate we will most of the time simply write set-theoretic solutions or solutions.
Rump has shown in [17] that the binary operation $x\cdot y=\lambda_{x}^{-1}(y)$ turns a solution $X$ into a particularly easy algebraic structure; a (non-degenerate) cycle set, which is defined as a set $X$ with a binary operation $(x,y)\mapsto x\cdot y$ such that the following axioms are satisfied:
@@
The cycle sets considered in this work will always be assumed to be of finite size and will be referred to as cycle sets, without the predicate non-degenerate. Cycle sets are in bijective correspondence with set-theoretic solutions to the Yang-Baxter equation. More precisely, given a cycle set $(X,\cdot)$ then $(X,r_{X})$ is a set-theoretic solution to the YBE, where $r_{X}:X^{2}\to X^{2}$ is given by
$r_{X}(x,y)=(\sigma_{x}^{-1}(y),\sigma_{x}^{-1}(y)\cdot x).$
Homomorphisms of cycle sets are defined in the obvious way. Consequently, this holds for all derived notions, such as isomorphisms. Note that isomorphisms of cycle sets translate to isomorphisms of set-theoretic solutions under the correspondence described above. As usual, if there is no danger of confusion, a cycle set $(X,\cdot)$ is simply referred to by its underlying set $X$.
Of particular importance for understanding general cycle sets is the class of indecomposable cycle sets. A cycle set $X$ is called indecomposable if there are no subsets $X_{1},X_{2}\neq\emptyset$ closed under the cycle set operation such that $X=X_{1}\cup X_{2}$. In a sense, they are elementary building blocks for cycle sets and turn out to be significantly more well-behaved. This can be seen from computations of Akgün, Mereb and Vendramin [1] that imply that the number of cycle sets of given cardinality grows in an aggressive manner, while the number of indecomposable cycle sets stays quite small. This indicates that there are strong obstructions for a cycle set to be indecomposable. Note that a cycle set $(X,\cdot)$ is indecomposable if and only if the corresponding set-theoretic solution $(X,r)$ is indecomposable in the sense that there is no partition $X=A\sqcup B$ ($A,B\neq\emptyset$) such that $r(A^{2})=A^{2}$ and $r(B^{2})=B^{2}$.
Indeed, up to isomorphism, there is only one indecomposable cycle set of size $p$, $p$ a prime number, that is given by $X=\mathbb{Z}_{p}$ with the operation $x\cdot y=y+1$ [9, Theorem 2.13]. Furthermore, Jedlička and Pilitowska [14] have recently classified all indecomposable cycle sets of multipermutation level 2 by group-theoretic means. In particular, this includes all indecomposable cycle sets of order $pq$ that are of finite multipermutation level, where $p,q$ are (not necessarily distinct) primes. Cedó and Okniński [6, Section 5] have described a class of simple cycle sets of size $p^{2}$, $p$ a prime. In [7], the same authors have recently proven that all indecomposable cycle sets of squarefree size are of finite multipermutation level.
In this article, we build on these results by providing a full classification of indecomposable cycle sets of size $p^{2}$, $p$ a prime. Using the correspondence between cycle sets and set-theoretic solutions, our final result, 5.1, reformulated in terms of set-theoretic solutions, will be as follows:
@@
Each non-degenerate, indecomposable, involutive set-theoretic solution $(X,r)$ to the Yang–Baxter equation of size $p^{2}$ for some prime $p$ is isomorphic to one of the following solutions:
@@
    $r\begin{pmatrix}(a,x)\\ (b,y)\end{pmatrix}=\begin{pmatrix}(b-1,y-\chi_{0}(b-1)S-\Phi(b-1-a))\\ (a+1,x+\chi_{0}(a)S+\Phi(a-b+1))\end{pmatrix}$
    where $\Phi:\mathbb{Z}_{p}\to\mathbb{Z}_{p}$ is a non-constant map with $\Phi(0)=0$, $S\in\mathbb{Z}_{p}$ and $\chi_{0}:\mathbb{Z}_{p}\to\mathbb{Z}_{p}$ with
@@
    The parameters $S,\Phi$ and $S^{\prime},\Phi^{\prime}$ define isomorphic solutions if and only if $S=S^{\prime}$ and $\alpha\Phi=\Phi^{\prime}$ for some $\alpha\in\mathbb{Z}_{p}^{\ast}$.
@@
    where $\Phi:\mathbb{Z}_{p}\to\mathbb{Z}_{p}$ is a non-constant map with $\Phi(x)=\Phi(-x)$ and $\alpha\in\mathbb{Z}_{p}^{\ast}$ is such that $\Phi(\alpha x)=\alpha\Phi(x)$.
    The parameters $\alpha,\Phi$ and $\alpha^{\prime},\Phi^{\prime}$ define isomorphic solutions if and only if $\alpha=\alpha^{\prime}$ and there is a $\beta\in\mathbb{Z}_{p}^{\ast}$ such that $\beta^{-1}\Phi(\beta x)=\Phi^{\prime}(x)$ for all $x\in\mathbb{Z}_{p}$.
Note that the classification of indecomposable retractable solutions of size $p^{2}$ follows from the theoretical classification of indecomposable solutions of multipermutation level $2$ by Jedlička and Pilitowska in [14]. In Section 2, we merely give them in a more explicit form building further upon their results, which also gives a more uniform presentation compared to the irretractable case.
@@ 1. Preliminaries
which is easily verified to be left distributive with respect to the addition. The set of fixed points of the $\lambda$-action is denoted by
$\operatorname{Fix}(A)=\{a\in A\colon b+a=b\circ a\text{ for all }b\in A\}=\{a\in A\mid b*a=0\text{ for all }b\in A\}.$
@@
Let $b\in B$. We have to show that $\lambda_{a}(b)\in B$ for all $a\in A$. Note that ${}^{a}b\in B$ because $B$ is normal in $A^{\circ}$, therefore $\lambda_{{}^{a}b}(a)\in a+B$ as $B^{\circ}$ acts trivially on $A^{+}/B^{+}$. Using 1.1 and the normality of $B^{+}$, we deduce that
@@
Let $A$ be a finite skew brace and $B$ a subbrace such that $p=|A|/|B|$ is equal to the least prime dividing $|A|$. Then $B$ is an ideal of $A$.
@@
$\operatorname{Fix}(A)$ is the set of fixed points under the action $\lambda:A^{\circ}\to\operatorname{Aut}(A^{+})$. It is well-known that an action of a $p$-group by automorphisms of a $p$-group always has non-trivial fixed points.
If we assume that $A^{\circ}$ is abelian, 1.2 implies that $\operatorname{Fix}(A)\subseteq\operatorname{Soc}(A),$ hence $\operatorname{Soc}(A)\neq 0$ as well. ∎
If $X$ is a set, we denote by $\operatorname{Sym}_{X}$ the symmetric group consisting of all permutations of $X$. If $n$ is a positive integer, we write $\operatorname{Sym}_{n}=\operatorname{Sym}_{\{1,\ldots,n\}}$.
Let $G$ be a group acting on a set $X$. A system of imprimitivity is a partition $X=\bigsqcup_{i\in I}X_{i}$ that is invariant under the group action in the sense that for any $g\in G$, $i\in I$, there is a $j\in I$ such that $g\cdot X_{i}=X_{j}$. The subsets $X_{i}$ are called blocks of the system. A system of imprimitivity is trivial if either $|I|=1$ or $|X_{i}|=1$ for all $i\in I$. Else it is non-trivial.
@@
Let $G$ be a non-abelian $p$-group with a transitive action on a set $X$ of size $p^{2}$. Then $G$ has at most one non-trivial system of imprimitivity.
@@
If there is more than one non-trivial system of imprimitivity, a theorem of Lucchini [16, Theorem 1] implies that $G\leqslant\operatorname{Sym}_{p}\times\operatorname{Sym}_{p}$. But if $G$ is a $p$-group, this forces $G$ to be isomorphic to a subgroup of $\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ and therefore to be abelian. ∎
Let $G,H$ be groups such that $G$ acts on the set $X$. Then the wreath product is defined as the semidirect product $G\wr_{X}H=G\ltimes H^{X}$ where $G$ acts on $H^{X}$ by $g\cdot(h_{x})_{x\in X}=(h_{g^{-1}\cdot x})_{x\in X}$. If the action of $G$ on $X$ is clear, we will generally suppress the subscript-$X$ and write $G\wr H$.
If additionally, $H$ acts on a set $Y$ then $G\wr H$ acts on $X\times Y$ by permutations of the form
@@
As a special case of [13, II, Satz 3.2], we know that every solvable transitive subgroup of $\operatorname{Sym}_{\mathbb{Z}_{p}}$ is conjugate to a subgroup of $\mathrm{AGL}(1,p)$. So we conclude that $G$ is conjugate to a subgroup of $\mathrm{AGL}(1,p)\wr\mathrm{AGL}(1,p)$. As $\mathbb{Z}_{p}\wr\mathbb{Z}_{p}$ is a $p$-Sylow subgroup of $\mathrm{AGL}(1,p)\wr\mathrm{AGL}(1,p)$ also the last part of the statement follows. ∎
@@
To a cycle set $(X,\cdot)$ we associate its structure group
@@
There is a unique way of defining an addition on $G(X)$ such that $x+y=x\circ\sigma_{x}(y)$ for $x,y\in X$ and $(G(X),+,\circ)$ is a brace, the structure brace of $X$. The canonical map $\iota:(X,\cdot)\to(G(X),\cdot):x\mapsto x$ is always an injective cycle set homomorphism, so we can identify $X$ with $\iota(X)$. Note that it follows immediately from the definition that $\iota(X)$ is a cycle base of $G(X)$.
@@
Recall that a cycle set $X$ is indecomposable if and only if $\mathcal{G}(X)$ acts transitively on $X$. The above group structure will also be denoted by $\mathcal{G}(X)^{\circ}$ and we obtain, as described in [5], a brace structure $(\mathcal{G}(X),+,\circ)$ by defining $\sigma_{x}^{-1}+\sigma_{y}^{-1}=\sigma_{x}^{-1}\circ\sigma^{-1}_{\sigma_{x}(y)}$ for $x,y\in X$. Note that this means that we obtain a surjective brace homomorphism $G(X)\twoheadrightarrow\mathcal{G}(X):x\mapsto\sigma_{x}^{-1}$, where the kernel is precisely $\operatorname{Soc}(G(X))$. As a result, the canonical map
@@
becomes a cycle set homomorphism. The image $\pi(X)$ is called the retraction of $X$, denoted $\mathrm{Ret}(X)$. If $\pi$ is injective, hence $(X,\cdot)$ is isomorphic to a sub-cycle set of $(\mathcal{G}(X),\cdot)$, then we say that $X$ is irretractable. When $\pi$ is not injective, $(X,\cdot)$ is retractable. Note that $\pi(x)=\pi(y)$ if and only if $\sigma_{x}=\sigma_{y}$. If moreover, we can obtain a one-element cycle set by iteratively repeating the retraction process on a cycle set $X$, we say that $X$ has finite multipermutation level. More precisely, if $k$ is the smallest value such that $|\operatorname{Ret}^{k}(X)|=1$, then we say that $X$ has multipermutation level $k$.
@@
Let $f:X\twoheadrightarrow Y$ be a surjective cycle set homomorphism between finite cycle sets. If $X$ is indecomposable, then $Y$ is also indecomposable and the fibres $f^{-1}(y)$ all have the same cardinality. In particular, $|Y|$ divides $|X|$.
2. Indecomposable retractable cycle sets of size $p^{2}$